OpenStudy (anonymous):

Verify the identity. cos 4x + cos 2x = 2 - 2 sin^2 2x - 2 sin^2 x

OpenStudy (superdavesuper):

OUCH...tis a tough one! first eqn u need is cos(2A) = cos^2(A) - sin^2(A) use that to change the cos 4x into cos^2 2x and sin^2 2x okay?

OpenStudy (anonymous):

Do you know Euler's identity?

OpenStudy (anonymous):

cos 4x + cos 2x = 2 - 2 sin^2 2x - 2 sin^2 x

OpenStudy (anonymous):

@superdavesuper

OpenStudy (raden):

you need this identity : cos(2x) = 1 - sin^2 (x) now can you identify for cos(4x) cos(4x) = cos(2 * (2x)) = ... ?

OpenStudy (raden):

oppps i mean cos(2x) = 1 - 2 sin^2 (x)

OpenStudy (superdavesuper):

cos 4x = cos 2(2x) and then use the general formula cos 2A = cos^2 (A) - sin^2 (A) to convert....